# Large Language Models Are Zero-Shot Time Series Forecasters

### v1

@@ C.2 Monash datasets
We evaluate on the top 16 datasets in Monash where the baseline methods perform the best, by sorting in decreasing order of the mean absolute error normalized by the MAE achieved by simply predicting the last observed value before the test series (a naive baseline), excluding datasets that contain missing values. The following datasets were selected: covid deaths, solar weekly, tourism monthly, australian electricity demand, pedestrian counts, traffic hourly, hospital, fred md, tourism yearly, tourism quarterly, us births,nn5 weekly,solar 10 minutes, traffic weekly, saugeenday, cif 2016.
Several of the baseline methods in the archive are shared with Darts, and all descriptions and code can be found in [Godahewa et al., 2021]. A few notable addition include
  • CatBoost: CatBoost [Prokhorenkova et al., 2018] is gradient-boosting framework for continuous or categorical data.
@@
We use the following hyperparameters for GPT-3: $\alpha=0.9,\beta=0,\mathrm{temperature}=0.7.$ To avoid exceeding the context window, we truncate the history to at most 500 most recent observations. For the baselines, we report their performance as presented in [Godahewa et al., 2021]. The normalized MAE values shown in Figure 4 (center) are obtained by normalizing by the lowest baseline MAE on each dataset before aggregating.
@@ C.3 Informer datasets
  • Informer: Informer [Zhou et al., 2021] is an efficient transformer model with sparse attention designed for long sequences.
  • Reformer: Reformer [Zhou et al., 2021] uses a locality-sensitive hashing mechanism to improve the memory use of attention.
  • Autoformer: Autoformer [Wu et al., 2021] is a model design for long time series that replaces standard attention with a mechanism in Fourier space.
  • FEDformer: Like Autoformer, FEDformer [Zhou et al., 2022] uses frequency-based decompositions to construct an efficient alternative to attention.
@@ C.5 Darts full probabilistic results
Figure 9: Median predictions of LLMTime (GPT-3) and NLLs from LLMTime (GPT-3 and LLaMA-2 70B) for every dataset within Darts [Herzen et al., 2022]. The shaded area shows the 10th to 90th quantiles of the distribution over samples. LLMTime consistently obtains better likelihood values than the baselines and often makes surprisingly accurate forecasts by effectively extrapolating trend and periodic components.
@@ C.8 Informer dataset visualizations
Figure 12: LLMTime (LLaMA-2 70B base model) median predictions on 4 randomly chosen series per Monash dataset.
@@ Appendix D Simplicity bias experiments
We generate data from the function $f(x)=\cos(x)+x$ and add Gaussian noise with zero mean and variance $0.05$. We fit symbolic expressions to the first 140 timesteps using PySR [Cranmer, 2023] with symbols ["$+$", $\cdot$", "-", "///", "$\sin$", "$\cos$", "$\exp$","square"] and $\mathrm{maxsize}=70,\mathrm{maxdepth}=10,\mathrm{population\_size}=50,\mathrm{loss}=\mathrm{abs(prediction-target)},\mathrm{model\_selection}=\mathrm{accuracy}$ and $\mathrm{niterations}=100$. The solutions are saved and ranked by complexity, which is simply the number of terms in the symbolic regression function. The five solutions shown in Figure 5 are

### v2

@@ C.2 Monash datasets
We evaluate on 19 datasets in Monash that satisfy two criteria:
  1. 1.
    The total number of individual series cannot be prohibitively large, so that the experiments can be run in time without access to an enormous cluster and without a gratuitous API expenses.
  2. 2.
    The length of the forecasting horizon cannot extend to a length that makes it impossible to fit both the forecast and the history into the context window of the language model.
When we applied these criteria, we obtained the following 19 datasets were selected: covid deaths, solar weekly, tourism monthly, australian electricity demand, pedestrian counts, traffic hourly, hospital, fred md, tourism yearly, tourism quarterly, us births, nn5 weekly, nn5 daily, traffic weekly, saugeenday, cif 2016, bitcoin, sunspot.
To aggregate across datasets, we normalized the mean absolute error by the MAE achieved by simply predicting the last observed value before the test series (a naive baseline). This normalization places high weight on datasets for which methods perform significantly better or worse than the naive predictor.
Several of the baseline methods in the archive are shared with Darts, and all descriptions and code can be found in [18]. A few notable addition include
  • CatBoost: CatBoost [39] is gradient-boosting framework for continuous or categorical data.
@@
We use the following hyperparameters for GPT-3: $\alpha=0.9,\beta=0,\mathrm{temperature}=0.7.$ To avoid exceeding the context window, we truncate the history to at most 500 most recent observations. For the baselines, we report their performance as presented in [18]. The normalized MAE values shown in Figure 4 (center) are obtained by normalizing by the lowest baseline MAE on each dataset before aggregating.
@@ C.3 Informer datasets
  • Informer: Informer [54] is an efficient transformer model with sparse attention designed for long sequences.
  • Reformer: Reformer [54] uses a locality-sensitive hashing mechanism to improve the memory use of attention.
  • Autoformer: Autoformer [49] is a model design for long time series that replaces standard attention with a mechanism in Fourier space.
  • FEDformer: Like Autoformer, FEDformer [55] uses frequency-based decompositions to construct an efficient alternative to attention.
@@ C.5 Darts full probabilistic results
Figure 9: Median predictions of LLMTime (GPT-3) and NLLs from LLMTime (GPT-3 and LLaMA-2 70B) for every dataset within Darts [23]. The shaded area shows the 10th to 90th quantiles of the distribution over samples. LLMTime consistently obtains better likelihood values than the baselines and often makes surprisingly accurate forecasts by effectively extrapolating trend and periodic components.
@@ C.8 Informer dataset visualizations
Figure 12: LLMTime (LLaMA-2 70B base model) median predictions on 4 randomly chosen series per Informer dataset.
@@ Appendix D Simplicity bias experiments
We generate data from the function $f(x)=\cos(x)+x$ and add Gaussian noise with zero mean and variance $0.05$. We fit symbolic expressions to the first 140 timesteps using PySR [14] with symbols ["$+$", $\cdot$", "-", "///", "$\sin$", "$\cos$", "$\exp$","square"] and $\mathrm{maxsize}=70,\mathrm{maxdepth}=10,\mathrm{population\_size}=50,\mathrm{loss}=\mathrm{abs(prediction-target)},\mathrm{model\_selection}=\mathrm{accuracy}$ and $\mathrm{niterations}=100$. The solutions are saved and ranked by complexity, which is simply the number of terms in the symbolic regression function. The five solutions shown in Figure 5 are